### v1

# DLoRA-TrOCR: Mixed Text Mode Optical Character Recognition Based On Transformer

Abstract
With the continuous development of OCR technology and the expansion of application fields, text recognition in complex scenes has become a key challenge.Factors such as multiple fonts, mixed scenes and complex layouts seriously affect the recognition accuracy of traditional OCR models. Although OCR models based on deep learning have performed well in specific fields or similar data sets in recent years, the generalization ability and robustness of the model are still a big challenge when facing complex environments with multiple scenes. Furthermore, training an OCR model from scratch or fine-tuning all parameters is very demanding on computing resources and inference time, which limits the flexibility of its application. This study focuses on a fundamental aspect of mixed text recognition in response to the challenges mentioned above, which involves effectively fine-tuning the pre-trained basic OCR model to demonstrate exceptional performance across various downstream tasks.To this end, we propose a parameter-efficient hybrid text recognition method based on pre-trained OCR Transformer, namely DLoRA-TrOCR. This method embeds DoRA into the image encoder and LoRA into the internal structure of the text decoder, enabling efficient parameter fine-tuning for downstream tasks.Experimental results show that compared to similar parameter adjustment methods, our model DLoRA-TrOCR has the smallest number of parameters and performs better. It can achieve state-of-the-art performance on complex scene data sets involving simultaneous recognition of mixed handwritten, printed and street view texts.
[…]
123
1 Introduction
Optical Character Recognition (OCR) is the technology of classifying optical patterns corresponding to letters, digits, or other characters in digital images[2]. It typically consists of two tasks: text detection and text recognition. Text detection involves detecting text regions from an image, while text recognition extracts and recognizes the text within those regions.
The main practical application of OCR is to read text contained in various types of images, commonly scanned handwritten or printed documents. Due to the different text media in digital images, the text formats for optical character recognition mainly include the following two types:
  • Formatted text: Standardized printed text following a given font and format, most commonly found in printed images.
  • Handwritten text: Text written by humans. This type of text is variable, irreproducible, and non-uniform.
Figure 1: Optical Character Recognition: Text Detection and Text Recognition
Initial testing and development of OCR methods are typically carried out on images containing machine-printed text[16]. Traditional OCR research has aimed to optimize the recognition of printed text, focusing on the text detection task. It was not until 20th century that the importance of OCR for recognizing handwritten text gained prominence[23]. English handwritten text, with its vast user base,varying personal handwriting styles, and differences in character shapes and sizes, poses higher demands on the text recognition task. Additionally, text recognition in more complex multi-scene backgrounds has received widespread attention. The ICDAR Robust Reading Competition released the first scene text detection and recognition benchmark, ICDAR 2003[1]. In the following two decades, competitions and test benchmarks related to scene recognition became more and more abundant and standard, and the methods involved, including deep neural networks, were increasingly used[14]. The increasingly complex scene background also poses challenges to the computer vision task of text detection[15].
Transformer[28] has developed rapidly in many fields since it was proposed, especially in general computer vision tasks, and the effectiveness of Transformer has been verified[21]. Recently, the OCR method of English text based on Transformer has made significant progress in text recognition performance in this field[12, 6]. It makes full use of the Transformer architecture to unify detection and text recognition into an end-to-end task.
Existing text recognition methods often default to training specialized models based on specific tasks. However, the performance of these models can significantly decline when processing images containing mixed text types[22]. Progress in OCR technology capable of handling multi-scene and multi-type scenarios has been hindered by the high computational demands of training parameter-heavy models on large datasets[21]. The recent rise of general large language models and the pre-training/fine-tuning paradigm in natural language processing has provided new directions for multi-scene text recognition in OCR.
Our research aims to extend the application of pre-training and fine-tuning to the OCR domain. To construct a versatile multi-scene text recognition model, we built a dataset containing a mixture of handwritten, printed, and scene text. We then leveraged the pre-trained Transformer OCR model TrOCR, utilizing its weights from training on synthetic handwritten and printed text, and fine-tuned it on our mixed dataset. Introducing parameter-efficient fine-tuning (PEFT) techniques, we adopted the DoRA and LoRA methods for the encoder and decoder, respectively. This significantly reduced the trainable parameters while effectively improving performance on the mixed dataset. Experiments demonstrated outstanding performance in handling complex scenarios with mixed handwritten, printed, and scene text using only a small number of trainable parameters.
Our key contributions are threefold:
  • We constructed a mixed dataset of English handwritten, printed, and complex scene text to cover diverse image formats.
  • Our work pioneers the application of a mixed-text OCR baseline model and a novel use of the PEFT method, addressing the inherent complexity of fine-tuning Transformer-based Vision-Language models.
  • We conducted comprehensive experiments on benchmark datasets for handwritten, printed, and scene text recognition tasks to validate the effectiveness and efficiency of our proposed method.
2 Related Work
2.1 OCR Framework.
OCR models usually adopt a pipeline structure composed of multiple independent modules, such as CNN (Convolutional Neural Network)[7] for feature extraction, LSTM (Long Short Term Memory)[35] for sequence modeling, etc. This compartmentalized design has many flaws, making it difficult for the model to fully utilize global information and long-range dependencies. In recent years, the Transformer model has achieved excellent performance in sequence modeling tasks due to its powerful self-attention mechanism and end-to-end architecture design[3].It mainly uses an encoder-decoder architecture to achieve from image to text in OCR tasks.
The SRN proposed by Yu et al[33], employs Transformer-encoder structure followed by the convolutional feature network to enhance visual features. In terms of decoder, the SRACN text detection algorithm proposed by Yang et al[31], uses Transformer-decoder to replace LSTM module, which once again verifies the efficiency and accuracy advantages of parallel training. The NRTR proposed by Sheng et al[25]. uses a complete Transformer architecture to encode and decode input images, and only uses a few simple convolutional layers for deep feature extraction to verify the effectiveness of the Transformer structure in text recognition. The TrOCR model proposed by Li et al. applies the pre-trained Transformer encoder-decoder framework to the OCR task. Unlike previous methods, TrOCR does not use CNN as the backbone network for image understanding, but directly uses the pre-trained image Transformer model as the encoder and the pre-trained text Transformer model as the decoder.
2.2 Parameter-Efficient Fine-Tuning Method.
The foundation model[36]has strong generalization ability, and more and more research attempts to use transfer learning to apply it to tasks in specific scenarios. However, the scale of modern deep learning and the requirements for computing scale make traditional transfer learning strategies gradually ineffective, and the need for efficient parameter adjustment becomes increasingly urgent[4]. Parameter-Efficient Fine-Tuning (PEFT)[19] aims to reduce the high cost of fine-tuning large models. This type of method achieves adaptation to downstream tasks by adjusting a relatively small subset of parameters, rather than adjusting all parameters of the model as in traditional methods. This strategy can significantly reduce computational cost and storage requirements compared to the total parameter amount.
Existing PEFT methods can be mainly divided into two categories: addition-based methods and reparameterization-based methods.Our research focuses on the second category of heavy-parameter-based methods, which are implemented by converting the adaptive parameters in the optimization process into a parametrically efficient form. This method was originally applied to natural language models. The representative of this type of method is Low-Rank Adaptation(LoRA)[8], which assumes that changes in weights during model adjustment have low intrinsic rank. Based on this assumption, LoRA proposes a low-order decomposition optimization approximation method for the changes in the original weight matrix in the self-attention module. LoRA can match the fine-tuned performance of the GLUE benchmark while adding almost no additional inference overhead. Subsequently, Liu et al. introduced the idea of weight normalization based on LoRA, decomposed the pre-training weight into two components, amplitude and direction, and used LoRA to update the direction component. This method, called DoRA (Weight-Decomposed Low-Rank Adaptation)[32], enhances the learning capabilities and training stability of LoRA on some tasks without introducing any additional inference overhead.
Figure 2: Transformer-based architecture, where the encoder-decoder model is a pre-trained image transformer as the encoder and a pre-trained text transformer as the decoder.Then, this model is pre-trained in two stages on a synthetic dataset of millions of handwritten and printed texts
3 Approach
Our model framework is based on the structure of TrOCR. In order to improve the generalization and robustness of the model when facing multi-scene and multi-task OCR recognition tasks, we introduce the PEFT framework, which reduces the amount of training parameters and improves inference efficiency while maintaining model accuracy. The specific measures are to adopt the DoRA method to improve the image encoder, and the LoRA method to improve the text decoder, which greatly reduces the amount of adjustment parameters and improves the overall performance of the encoder-decoder module.
3.1 Transformer-based OCR
Architecture of OCR Transformer.
The Transformer architecture we adopt includes an image Transformer for extracting visual features and a text Transformer for language modeling. The encoder aims to obtain a representation of image patches, while the decoder generates a sequence of wordpieces guided by visual features and previous predictions. Our model architecture still uses the original Transformer encoder-decoder structure. The encoder aims to obtain a representation of image patches, while the decoder generates a sequence of wordpieces guided by visual features and previous predictions.
Image Encoder.
The image encoder we adopt is based on the ViT model, which receives the input image $x_{\mathrm{img}}\in\mathbb{R}^{3\times H_{0}\times W_{0}}$ and resizes it to a fixed size $(H,W)$. Then, the input image is divided into $N$ image blocks of fixed size (P, P), tiled into vectors, and linearly projected to D-dimensional vectors. To preserve the position information of each pixel in the encoded image, absolute position embeddings are added to the embedding vector of each image patch. Next, the sequence of vectors is fed into a series of identical Transformer encoder layers. Each Transformer encoder layer contains a self-attention layer and a feed-forward layer. After each layer, a normalization layer and residual connections are applied. The self-attention mechanism calculates the similarity between each query and all keys, and uses the similarity as a weight to weight all values, and finally obtains the output. The feedforward layer usually consists of two fully connected layers and uses an activation function (such as ReLU and Swish[13]) for nonlinearization.
$\operatorname{Self-Att}(Q,K,V)$ $=V\operatorname{softmax}\big{(}\frac{K^{\top}Q}{\sqrt{D_{k}}}\big{)},$
$\operatorname{MultiHead}(Q,K,V)$ $=W_{o}[\operatorname{head}_{1};\cdots;\operatorname{head}_{h}],$
$\operatorname{head}_{i}$ $=\operatorname{Self-Att}(Q_{i},K_{i},V_{i}),$
$\forall i\in\{1,\cdots,h\},\quad Q_{i}$ $=W_{q}^{i}Q,K_{i}=W_{k}^{i}K,V_{i}=W_{v}^{i}V.$
where $Q$, $K$ and $V$ are the query, key and value matrices respectively, and $D$ is the dimension of the input vector. The self-attention layer uses linear projection to obtain query, key and value from the input sequence. Compared to a single attention function, each attention module utilizes multiple heads with different parameters, allowing the model to collect information from different representation subspaces.
Text Decoder.
The text decoder is based on the RoBERT model. Different from the image encoder, the decoder inserts an encoder-decoder attention layer after the multi-head self-attention layer to allocate different attention on the output of the encoder. In the encoder-decoder attention layer, the key and value come from the encoder output, while the query comes from the decoder input. The encoder-decoder attention layer weights all values by calculating the similarity between the query and the key and using the similarity as a weight. Although the part of the text decoder is initialized by the encoder-only model, the autoregressive constraint is still introduced during training, that is, the generated partial output and the whole input sequence can only be relied on to predict the next word, and the subsequent target word cannot be directly accessed. For the output sequence $y=(y_{1},y_{2},...,y_{n})$, given the conditional distribution $P(y\mid x)$of the input sequence $x$, the decoder generates the output sequence word by word in an autoregressive way, then:
$P(y\mid x)=\prod_{t=1}^{n}P(y_{t}\mid y_{<t},x)$
Where $y_{<t}$ denotes all generated words up to time $t$. The model is trained by computing the negative log-likelihood loss as follows,
$\mathcal{L}=-\sum_{i}\log P(y_{i}\mid x_{i})$
in the inference phase, for each time instant $t$, the decoder computes the probability of all possible words $w$ and applies the softmax function to obtain the normalized probability distribution: :
$y_{tw}$ $=Proj(Emb(w_{t}))$
$\text{Softmax}(y_{tw})$ $=\frac{e^{y_{tw}}}{\sum_{v=1}^{V}e^{y_{tv}}}\quad\text{for }w=1,2,...,V$
Here $V$ denotes the vocabulary size, $Emb(w_{t})$ denotes the embedding vector of the word $w_{t}$, and $Proj(\cdot)$ is the linear transformation that maps the embedding vector to the hidden state space.
Figure 3: Framework diagram of LoRA and DoRA methods.It is intuitive to see that after DoRA decomposes the pre-trained weights into directions and magnitudes, it updates the increments using LoRA in directions before computing with magnitudes
3.2 Parameter-efficient Fine-tuning
Although Transformer models perform well in task-specific OCR, they often contain billions of parameters, making fine-tuning and deploying these huge models in complex scenarios very resource-intensive. To overcome this challenge, we adopt PEFT Method, which can significantly reduce the number of parameters required for fine-tuning while maintaining model performance. Specifically, we used two PEFT methods: DoRA for fine-tuning the image encoder and LoRA for fine-tuning the text decoder.
LoRA
LoRA is a lightweight fine-tuning technology that achieves efficient fine-tuning by adding two small matrices to the key layers of the pre-trained model. Given the original weight matrix $W^{0}\in\mathbb{R}^{m\times n}$ of a pre-trained model layer, LoRA reparameterizes it as:
$W=W_{0}+BA$
where $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ are two small matrices, called LoRA weights, which are parameters that need to be learned during the fine-tuning process. $r$ is a hyperparameter that controls the rank of LoRA weight. When $r\ll\min(m,n)$, the number of parameters required by the LoRA weight will be much smaller than the original model weight matrix. By adding this low-rank update, LoRA enables efficient fine-tuning of pre-trained models. During the inference phase, the fine-tuned LoRA weights are combined with the original model weights to obtain the final inference weights. Since only two small matrices are stored, LoRA greatly reduces the additional parameter storage space required for fine-tuning. In our text decoder, we apply LoRA to both the query, key, and value projection matrices of the self-attention module, as well as the two linear layers of the feedforward network. This parameter-efficient fine-tuning method allows us to significantly reduce the number of parameters required to fine-tune the text decoder while maintaining performance. Specifically, LoRA’s weight matrix is used in Transformer block to reshape the multi-head attention mechanism, the incremental update of which can be expressed as
$Q=(W_{q}+B_{q}A_{q})\mathbf{x}$
$K=(W_{k}+B_{k}A_{k})\mathbf{x}$
$V=(W_{v}+B_{v}A_{v})\mathbf{x}.$
DoRA
DoRA builds upon LoRA technology. During fine-tuning, DoRA decomposes the pre-trained weight into two components: a magnitude vector $\mathbf{u}\in\mathbb{R}^{1\times n}$ vector and a direction matrix $V\in\mathbb{R}^{m\times n}$ matrix.It then leverages LoRA for efficient directional updates. This allows the incremental update to be expressed as:
$\Delta W=\mathbf{u}\frac{V+\Delta V}{\|V+\Delta V\|_{c}}=\|W_{0}\|_{c}\frac{W_{0}+BA}{\|W_{0}+BA\|_{c}}$
Notably, the magnitude vector continues to utilize the original $\|W_{0}\|_{c}$ obtained from the pre-trained weight decomposition, where $\|\cdot\|_{c}$ denotes column-wise normalization.
DoRA specifically targets the weight matrix of the multi-head attention mechanism within each Transformer encoder block. The updates of query, key and value in the self-attention mechanism are as follows:
$Q=(W_{q}+\|W_{q}\|_{c}\frac{W_{q}+B_{q}A_{q}}{\|W_{q}+B_{q}A_{q}\|_{c}})\mathbf{x}$
$K=(W_{k}+\|W_{k}\|_{c}\frac{W_{k}+B_{k}A_{k}}{\|W_{k}+B_{k}A_{k}\|_{c}})\mathbf{x}$
$V=(W_{v}+\|W_{v}\|_{c}\frac{W_{v}+B_{v}A_{v}}{\|W_{v}+B_{v}A_{v}\|_{c}})\mathbf{x}.$
An update of $\Delta W$ essentially corresponds to an update of the low-rank matrices $A$ and $B$ within $\Delta V$.From a gradient perspective, we have:
$\nabla_{A}\mathcal{L}=B^{\top}\frac{\partial\mathcal{L}}{\partial V^{\prime}},\nabla_{B}\mathcal{L}=\frac{\partial\mathcal{L}}{\partial V^{\prime}}A^{\top}$
where
$\frac{\partial{\mathcal{L}}}{\partial{V^{\prime}}}=\frac{\mathbf{u}}{\|V^{\prime}\|_{c}}\left(I-\frac{V^{\prime}V^{\prime\top}}{\|V\|_{c}^{2}}\right)\frac{\partial{\mathcal{L}}}{\partial{W}}$
While DoRA introduces an additional computational term $\frac{\mathbf{u}}{\|V^{\prime}\|_{c}}\left(I-\frac{V^{\prime}V^{\prime\top}}{\|V\|_{c}^{2}}\right)$ compared to LoRA,the impact on training efficiency is minimal.This slight overhead is a worthwhile trade-off for the gain in accuracy achieved by DoRA.
Table 1: Optimization of computational resources and model parameters using LoRA and DoRA methods, which load the pre-trained weights of the model using fp16, can actually be loaded with less precision in practice, which means that a larger batch size can be used on the same memory size, resulting in faster training.
Model PEFT Method GPU Memory Cost Tranable Params Total Params # Params
TrOCR-base - 23.28 333.9M 333.9M 100
TrOCR-base LoRA 15.01 2.0M 335.9M 0.585
TrOCR-base DoRA 16.44 2.1M 335.9M 0.614
TrOCR-base DLoRA 15.30 2.0M 335.9M 0.594
4 Experiments
4.1 Experiment Setup
4.1.1 Dataset.
Handwritten dataset:
From the IAM dataset[20], which contains handwritten English text and is one of the most commonly used datasets in the field of handwritten text recognition. We use the partitioning of the Aachen dataset: the training set contains 6,482 lines of text in 747 forms, the validation set contains 966 lines of text in 115 forms, and the test set contains 2,915 lines of text in 336 forms.
Printed dataset:
From the task2 of SROIE dataset[9], which focuses on text recognition in receipt images. The SROIE training set and test set contain 626 and 361 receipt images respectively. After cropping and removing duplicate images, the training set contains 10,682 lines of text, and the test set contains 6,897 lines of text.
Street View Text datasets:
From multiple public benchmarks, including IIIT5K-Words (IIIT5K)[17], Street View Text(SVT)[29], ICDAR 2013 (IC13)[10], ICDAR 2015(IC15)[11], Street View Text-Perspective (SVTP)[18], and CUTE80 (CUTE)[24]. These datasets contain various scene text images with noise such as blur, occlusion, or low resolution. The training set contains 7573 lines of text, and the test set contains 11435 lines of text. The specific data set situation is shown in the Table 2. To achieve better generalization on multi-scenario tasks, we adopt hybrid measures on the dataset. Compared with the circular sampling method [27]that may lead to errors in uneven data distribution, we adopt a direct and balanced data integration strategy to directly mix the data sets of the above-mentioned multiple related tasks, eliminating the possibility of a single task dominating the training process. nature, promoting the comprehensive learning of the model.
Table 2: Handwritten, printed, and street view text datasets.Dataset Size represents the number of characters.The numbers in parentheses represent the number of lines of text that have been cropped.
Dataset type Dataset Train set size Test set size Crop
Handwritten dataset IAM 747 (6842) 336 (2915)
Printed dataset SROIE Task2 626 (10682) 361 (6897)
STR benchmarks IC15 4468 3888
IC13 848 2967
III5K 2000 3000
SVT 87(257) 249(647)
SVTP - 645
CUTE - 288
The datasets used in our subsequent comparative experiments, combination experiments, and ablation experiments are all trained and tested on the above-mentioned handwritten text dataset, printed text dataset, scene text dataset, and our mixed text dataset.
4.1.2 Evaluation Metric.
We considered two evaluation metrics that are widely used in OCR systems, Character Error Rate(CER) and F1 score. CER is designed to be evaluated at the character level, which is case-sensitive, while the F1 score is evaluated at the word level, which is case-insensitive. CER is defined as the edit distance between the predicted text and the ground truth. CER can be calculated as:
$\mathrm{CER}=\frac{S+I+D}{N}$
where $S$ is the number of substitutions, $I$ is the number of insertions, $D$ is the number of deletions, and $N$ is the total number of characters in the real text.
WER and CER are similar benchmark scores. The main difference is that the smallest unit of text in the calculation method is changed from characters to words, so that WAR can be calculated by $1-WER$. We mainly use it in the evaluation of the STR data set.
The F1 score is the harmonic mean of precision and recall, used to evaluate the accuracy of a model, which is defined as follows:
[…]
Precision=Correct matchesThe number of detected $\mathrm{Precision}=\frac{\text{Correct matches}}{\text{The numberof detected words}}$
Recall=Correct matchesThe number of ground truth $\mathrm{Recall}=\frac{\text{Correct matches}}{\text{The number ofground truth words}}$
4.1.3 Implementation Details.
In the model configuration, we use the official pre-training weights of the TrOCR-Base model, and use the AdamW optimizer to perform training for 20 epochs and a batch size of 16. The initial learning rate of full-parameter Fine-Tuning is 1e-5, and the initial learning rate of PEFT method is 5e-5. Our comparative model experiments are all based on PyTorch, and all experiments are performed on an Nvidia RTX 4090 GPU with 24 GB of memory. In order to accurately evaluate the model performance and ensure the rationality of the training, validation and testing processes, we divided the mixed data set into a training set and a validation set according to the classic 9:1 ratio. In addition, we also retain an independent test set before mixing to ensure that the evaluation of the final performance of the model is completely independent of the training process.
4.2 Experiment Result
4.2.1 Performance.
We conducted a detailed comparative analysis between the proposed method and a variety of current state-of-the-art OCR technologies. The evaluation covers seven benchmark datasets, and adopts F1 and CER as core evaluation indicators. Detailed results are detailed in Table3.
The evaluation covers the above-mentioned benchmark data sets, and uses F1 and CER as core evaluation indicators as comparison standards. See Table 3 for detailed results. The results show that our model shows superior performance over contrasting methods on most evaluation datasets and corresponding metrics, reaching an accuracy of 84.63% on the hybrid dataset we constructed. This clearly confirms the effectiveness and advantages of combining pre-trained models with efficient fine-tuning of parameters within a multi-task text framework. Although the optimal performance is not achieved on the specific street view text recognition task, we believe this is because the optimal parameter weights of the TrOCR architecture are not disclosed, and we can only adopt the non-optimal weight parameters of the TrOCR architecture, which means that we There is still space for improvement in our model performance.
Table 3: Different models were simultaneously evaluated in IAM,SROIE, STR benchmark and Mix Dataset average CER and average F1 score.
Model IAM SROIE STR BENCHMARKS MIX DATASETS
F1 CER ACC F1 CER ACC ACC ACC
CRNN(Shi,Bai,and Yao 2016)[26] 66.30 10.8 65.87 80.92 - 84.47 79.55 76.63
SRN(Yu et al.2020)[34] 77.36 9.11 72.12 87.33 - 88.66 82.71 81.16
VisionLAN(Wang et al.2021)[30] 75.68 7.99 77.65 92.22 - 90.07 81.69 83.11
ABINet(Fang et al.2021)[5] 78.66 8.87 79.88 90.75 - 89.22 83.47 84.19
Ours 80.73 7.56 80.54 92.93 1.42 91.05 82.30 84.63
4.2.2 Result analysis based on PEFT method.
During the experiment, we compared the performance of efficient fine-tuning methods for different parameters, and the results are shown in Table 4.
Table 4: Comparative experimental results for full parameter fine-tuning,encoder-Decoder using LoRA or DoRA,encoder using DoRA and decoder using LoRA
Model IAM SROIE STR BENCHMARS
F1 CER F1 CER F1 CER
Fine-tuning 76.42 10.47 91.48 1.84 82.57 7.29
LoRA 80.71 7.47 92.74 1.47 83.31 6.88
DoRA 80.60 7.57 92.60 1.48 83.26 7.04
DLoRA 80.73 7.56 92.92 1.42 83.45 6.70
Compared with other methods, the excellent performance of DLoRA shows great potential for improving the accuracy of multi-task text OCR. It shows that the design of using different PEFT methods for the encoder part and decoder part of the Vision-Language encoder-decoder architecture is effective.
4.3 Ablation Study.
In order to prove the optimal combination of functions of the PEFT method, we designed a set of ablation experiments, and the result is shown in Table 5 and figure4. When encoder uses DoRA method and decoder uses LoRA method, the performance is significantly improved, with CER score of 5.42 and F1 score of 85.07, showing the best overall performance. The results of ablation experiments show that:
  • In OCR task, DoRA method is applied to a separate image encoder, which shows a slight improvement compared to LoRA method. We believe this is because DoRA can effectively suppress the influence of background noise on feature extraction through weight normalization, and enhance the robustness of the model to image input.
  • Full-parameter fine-tuning of only the image encoder is even better than fine-tuning of the image encoder and text decoder at the same time, because the pre-trained text decoder already has strong generalization in text decoding. performance, full-parameter fine-tuning on our relatively small mixed data set may destroy the original generalization.
  • Applying PEFT technology to encoders is far better than applying it to decoders. We believe the main reasons are:
    • 1.
      Diversity of information: Image encoders need to process diverse information from different types of images, while text decoders process relatively homogeneous text. Therefore, it may not be obvious whether the text decoder is processed or not, but it will be very important whether the image encoder is processed or not.
    • 2.
      Purpose of feature extraction Difference: The main purpose of the image encoder is to extract key features related to the text content from the image, while the text decoder predicts the next character according to the feature information transmitted by the image encoder combined with its previous prediction results. Therefore, the quality of the features extracted by the encoder has a more critical and direct impact on the performance of the decoder.
Table 5: Ablation experiments were performed in encoders and decoders using LoRA and DoRA methods and full parametric fine-tuning.
Encoder Decoder Avg. CER Avg. F1
- LoRA 14.22 72.36
- DoRA 14.17 72.45
- Fine-Tune 14.85 71.61
LoRA - 6.20 82.89
DoRA - 6.09 82.89
Fine-Tune - 5.35 84.52
LoRA LoRA 5.58 84.80
DoRA DoRA 5.56 84.89
LoRA DoRA 5.78 84.22
DoRA LoRA 5.42 85.07
Fine-Tune Fine-Tune 6.97 82.16
5 Conclusion
In this paper, we introduce the DoRA and LoRA methods of PEFT technology into the TrOCR model structure to enhance OCR’s recognition accuracy and multi-data set generalization ability in mixed text complex scenes. We use DoRA to decompose pre-training weights in the Encoder of Transformer structure, and only carry out LoRA updates for directional components to stably process multi-text images. The Decoder uses LoRA to update the attention weight and improve the decoding ability of mixed text features. In order to test the robustness of the model, we also constructed a complex scene dataset covering handwritten, printed, and scene view text, and verified the superiority of our model in terms of accuracy and number of parameters through a large number of comparative tests and ablation experiments. Our research makes new breakthroughs in hybrid text OCR and provides ideas on how to effectively adjust parameters in existing specific models. This lays the foundation for future research into more advanced, task-specific OCR methods.
Figure 4: Impact of DoRA and LoRA methods on the Encoder and Decoder modules respectively on our model performance.